# Quantum chaos in $\mathcal{PT}$-symmetric quantum systems

### v1

@@ I   Introduction
The out-of-time-order correlator (OTOC) [17, 18, 19, 16] is another possible measure of quantum chaos that may be used to identify an analogue of the Lyapunov exponent, providing a connection with classical chaos, e.g., via the butterfly effect. Previously, OTOCs have been calculated in the context of information scrambling [20, 21, 22, 23], quantum butterfly effects [24], many-body localization [25], “fast scrambling” [26, 27, 28], dynamical and topological phase transitions [29, 30, 31, 32], and open quantum systems [33, 34]. Recently, the experimental implementation of many-body time-reversal protocols in atomic quantum systems has attracted attention for its potential to measure OTOCs experimentally  [35, 36, 37, 38] leading to several specific experimental proposals to measure OTOCs and also the first experimental results [39, 15, 40, 41]
The OTOC has been calculated for the Hermitian kicked rotor [42]. It has been observed that at initial times, the OTOC exhibits an exponential growth in the chaotic regime, which is absent in the integrable regime. Therefore, it has been possible to extract a quantum equivalent of the classical Lyapunov exponent, at least in a semi-classical limit.
@@
Since the PTKR model is non-Hermitian and may have complex eigenvalues, we must use generalizations of the previously discussed diagnostics of quantum chaos [49] that apply to complex eigenvalues. For example, we study the complex energy level spacing ratio (CLSR) [50, 51] instead of the conventional real energy level spacing ratio (RLSR) [9]. Additionally, we calculate a suitably defined OTOC for non-Hermitian systems. [52].
Very recently, the time evolution of the OTOC has been studied for the PTKR model and its classical counterpart in the vicinity of a $\mathcal{PT}$ symmetry-breaking transition in parameter space. [53, 52, 54]. Our study extends this study to include the entire phase diagram of the PTKR model. It provides critical insights into the interplay of the $\mathcal{PT}$ symmetry-breaking transition and the transition from integrability to chaos.
The rest of the paper is structured as follows. In section II, we introduce the PTKR model. In section III, we discuss the two diagnostics of quantum chaos – complex level spacing ratio and out-of-ordered correlator. The main results based on these diagnostics are presented in section IV. Finally, we conclude in section V discussing our work’s implications and future directions.
II   Model
@@
$\mathcal{P}\hat{\theta}\mathcal{P}^{-1}-\hat{\theta}\,,\quad\quad\mathcal{P}\hat{p}\mathcal{P}^{-1}=-\hat{p}$ (3)
@@
III   Methods
III.1   Energy Level Spacing Ratio
@@
$\xi_{K}=\dfrac{z_{K}^{\text{NN}}-z_{K}}{z_{K}^{\text{NNN}}-z_{K}}=r_{K}\exp{\iota\theta_{K}}$ ) (7)
The complex level spacing ratio (CLSR) can be defined in two ways: One is by averaging over the magnitudes of $\xi$ to obtain $\langle r\rangle$ (derived from the set of $r_{K}$ values), and the other is averaging over the angular distribution to obtain $-\langle\cos{\theta}\rangle$ ) ⟩ (derived from the set of $\theta_{K}$ values). Only $\langle r\rangle$ can be defined in the limit when the spectrum becomes real. However, in this limit, Eqn. 7 does not yield the standard real level spacing ratio (RLSR) for a real spectrum. We, henceforth, refer only to $\langle r\rangle$ as the complex level spacing ratio (CLSR). Calculations of $-\langle\cos{\theta}\rangle$ ) ⟩ can be found in the supplementary material. The universality classes for random matrices that describe Hermitian matrices in terms of Gaussian Ensembles have suitable generalization to complex matrices [49]. We calculate the CLSR for these universality classes, one of which the PTKR model belongs.
There are two computational techniques to improve the statistics in our calculation. The first is to replace $m$ in 1 by $m+\Delta m_{p}$, where $\Delta m_{p}$ is a small random number selected independently for each $p$. This ensures that no underlying symmetries leading to unwanted degeneracies occur. This is important as degeneracies must be avoided when calculating level spacing ratios. The second is to average over a small window of values of the kicking strength $K$. This is done to avoid any accidental resonance in the $K$ values. Additionally, it is computationally inexpensive to average over different values of $K$ for the same system size rather than increase it.
III.2   Out-of-time-order correlator
@@
We choose $p_{0}=\hbar_{\text{eff}}k_{0}\in[-\pi,\pi]$ and $\sigma=4$. Numerically, $|\Psi\rangle$ is represented in a finite basis of eigenstates $|k\rangle$, $k\in[-N;N-1]$. The Hamiltonian is expressed as a matrix in this basis of $|k\rangle$ states, and all calculations are also performed w.r.t this basis. In both, the calculation of the OTOC and CLSR, an “effective value” of the Planck constant $\hbar_{\text{eff}}$ has been employed as a parameter to tune the amount of discreetness of the momentum operator [42]. This allows one to make a connection with the classical model by taking the limit $\hbar_{\text{eff}}\rightarrow 0$. We find no qualitative difference in our main results by changing $\hbar_{\text{eff}}$ and thus report them only for specific values of $\hbar_{\text{eff}}$ as mentioned in the appropriate parts of this paper.
Figure 1: The mean CLSR and RLSR as a function of the kicking strength $K$, calculated for $N=8005$ with $\hbar=0.2$. Horizontal lines represent analytical values of RLSR for Poisson and GOE statistics. The figure shows that the CLSR is as good an indicator of the transition from integrability to chaos as the more commonly employed RLSR.
Figure 2: The mean CLSR and RLSR as a function of the non-Hermiticity parameter $\lambda$, calculated for $N=4095$ with $\hbar=0.2$ and $K=0.15$. Horizontal lines represent the values of CLSR for the Poisson and GOE distributions. The above figure resembles the transition observed in  2. However, it is obtained by tuning $\lambda$ instead of $K$. This suggests that the system can transition from a $\mathcal{PT}$-symmetric integrable phase to a $\mathcal{PT}$-symmetric chaotic phase with increasing non-Hermiticity while the value of $K$ is kept constant.
@@
Figure 3: Left: The mean CLSR for varying values of the kicking strength $K$ and the non-Hermiticity $\lambda$, computed for $N=6005$. The dotted lines T1 and T2 represents the transitions across the $\mathcal{PT}$-symmetric integrable pahse to $\mathcal{PT}$ symmetry broken chaotic phase and the $\mathcal{PT}$-symmetric chaotic phase to $\mathcal{PT}$ symmetry broken chaotic phase, respectively. Line T* is the Hermitian transition from the integrable to chaotic phase. T3 marks a transition similar to T* in which $K$ is held constant and $\lambda$ varied. Center: The mean CLSR across the transition Main panel: T1 Inset T2, calculated for $N=4095$. The horizontal lines show the standard values of the mean CLSR for the GOE and Poisson distributions. Right: The absolute value of the maximum imaginary part of an energy eigenvalue, $\alpha$, across the transition Main panel: T1 Inset T2, calculated for $N=4095$. It can be seen that while $\alpha$ shows an abrupt change along T2, it seems to increase smoothly along T1. The CLSR on the other hand, shows an abrupt transition along both T1 and T2.
IV   Results
We now present the results of the calculations discussed in Sec. III for the PTKR model.
IV.1   CLSR
@@
For the Hermitian case ($\lambda=0$), the mean RLSR can be used to identify the transition from integrability to chaos. In what follows, we show that the CLSR can also be used for the same purpose and takes on a specific characteristic value in the two regimes.
@@
In Fig. 2, we show the mean RLSR as well as the mean CLSR with varying values of the kicking strength $K$, for $\hbar=0.2$ and system-size $N=8005$. We notice that the transition points are reasonably independent of the value of $N$. Thus, in what follows, the system size is chosen to be large enough so that all quantities are well converged. We find that both the CLSR and RLSR display a transition at the same value of $K$. However, the values in the two regimes are different. Furthermore, the values of the CLSR and RLSR that are obtained by the PTKR model in the chaotic limit of the Hermitian case match with those obtained from averaging over purely random matrices in the Gaussian Orthogonal Ensemble (GOE). The transition in the Hermitian case has been highlighted in the color plot as T* for aid in comparison.
@@
We now consider the non-Hermitian case ($\lambda\neq 0$), which is of particular interest in this study. In Fig. 3, we show the behavior of the mean CLSR for varying values of kicking strength $K$ and the non-Hermiticity $\lambda$, computed for $\hbar_{\text{eff}}=0.2$, and system-size $N=6005$. We define $\alpha=\text{max}\{\text{Im}(E)\}$ which only becomes non-zero when the spectrum starts to possess imaginary eigenvalues of the energy or, equivalently, when $\mathcal{PT}$-symmetry breaks (see supplementary material). We find that the phase diagram derived from the mean CLSR consists of three regimes: a $\mathcal{PT}$-symmetric integrable phase, a $\mathcal{PT}$-symmetric chaotic phase, and $\mathcal{PT}$-symmetry broken chaotic phase.
@@
There are three possible transitions that can occur: 1) From the $\mathcal{PT}$-symmetric integrable phase to the $\mathcal{PT}$-symmetry broken chaotic phase, which we label T1, 2) from the $\mathcal{PT}$-symmetric chaotic phase to the $\mathcal{PT}$-symmetry broken chaotic phase, which we label T2 and 3) from the $\mathcal{PT}$-symmetric integrable phase to $\mathcal{PT}$-symmetric chaotic phase, which we label T3. These are shown in Fig. 3. T3 is similar to the transition seen in the Hermitian case. This is because, along T3, we see that the CLSR goes from 0.50 to 0.57, which is exactly what is observed when going from the integrable to the chaotic regime for $\lambda=0$. For transition T2, we see a corresponding abrupt change in the value of $\alpha$ across the transition from the $\mathcal{PT}$-symmetric integrable regime to the $\mathcal{PT}$-symmetry broken chaotic regime. It can be seen that $\mathcal{PT}$-symmetry breaking at the transition T1 does not occur abruptly as can be expected for a system with finite Hilbert space dimension (See Fig. 3). Therefore, the exact position of the transition point cannot be determined with great precision; however, the CLSR does show a clear transition. We define a threshold value of $\alpha$ above which the $\mathcal{PT}$-symmetry is assumed to be broken. The threshold, keeping in mind numerical errors, is taken to be $\alpha\leq 10^{-10}$ for unbroken $\mathcal{PT}$-symmetry. We observe that the CLSR takes on the same value of 0.57 in the $\mathcal{PT}$-symmetric chaotic regime as it does in the chaotic regime of the Hermitian model. However, we find that in the $\mathcal{PT}$-symmetric chaotic regime, the value of $\alpha$ can be greater than the threshold of $10^{-10}$, demonstrating that the highest imaginary value of the eigenenergies is not a particularly good diagnostic to detect the absence of $\mathcal{PT}$-symmetry.
@@
It is interesting to note that the transition from the $\mathcal{PT}$-symmetric integrable phase to the $\mathcal{PT}$-symmetric chaotic interestingly also occurs across T3. The hermitian case $\lambda=0$ which is a transition from $\mathcal{PT}$- symmetric integrable to $\mathcal{PT}$-symmetric chaotic and is seen at T* and hence a similar transition can is expected when one varies $K$ for small enough $\lambda$ but not zero, which is also observed in the color plot. However, T3 involves varying $\lambda$ to get such a transition, not $K$. This is somewhat surprising as it seems to lack any Hermitian counterpart. One is able to induce chaos, as seen from the measured CLSR values, by increasing $\lambda$ whilst holding $K$ to the traditional integrable value 2. In the Hermitian model, the kicking strength must be varied to induce an analogous transition. Thus, the nature of this transition is intriguing and requires further study, which we defer to future work.
@@
Figure 4: The OTOC $C(t)$ and state norm $\mathcal{N}(t)$ vs $t$ across the transition along Left T1 and Right T2 shown in Fig. 3, calculated for system size $N=2^{14}$. For T1 (Left), the values $(\lambda,K)=\{(0.001,0.1),(0.002,0.456),(0.003,1.135)\}$ correspond to the $\mathcal{PT}$-symmetry unbroken integrable phase, while the values $(\lambda,K)=\{(0.006,5.734),(0.01,15.0)\}$ correspond to the $\mathcal{PT}$-symmetry broken chaotic phase. For T2 (right), the values $(\lambda,K)=\{(10^{-5},30),(10^{-4},30),(10^{-3},30)\}$ correspond to the $\mathcal{PT}$-symmetry unbroken chaotic phase, while the values $(\lambda,K)=\{(0.005,30),(0.001,30.0)\}$ correspond to the $\mathcal{PT}$-symmetry broken chaotic phase. In both cases, the $\mathcal{PT}$-symmetric broken phase dislays a late-time exponential growth of the OTOC and the state norm.
IV.2   OTOC

### v2

@@ 1 Introduction
The out-of-time-order correlator (OTOC) [17, 18, 19, 16] is another possible measure of quantum chaos that may be used to identify an analogue of the Lyapunov exponent, providing a connection with classical chaos, e.g., via the butterfly effect. Previously, OTOCs have been calculated in the context of information scrambling [20, 21, 22, 23], quantum butterfly effects [24], many-body localization [25], “fast scrambling” [26, 27, 28], dynamical and topological phase transitions [29, 30, 31, 32], and open quantum systems [33, 34]. Recently, the experimental implementation of many-body time-reversal protocols in atomic quantum systems has attracted attention for its potential to measure OTOCs experimentally  [35, 36, 37, 38] leading to several specific experimental proposals to measure OTOCs and also the first experimental demonstrations [39, 15, 40, 41]
The OTOC has been calculated for the Hermitian kicked rotor [42]. It has been observed that at initial times, the OTOC exhibits exponential growth in the chaotic regime, which is absent in the integrable regime. Therefore, it has been possible to extract a quantum equivalent of the classical Lyapunov exponent, at least in a semi-classical limit.
@@
Since the PTKR model is non-Hermitian and may have complex eigenvalues, we must use generalizations of the previously discussed diagnostics of quantum chaos [49] that apply to complex eigenvalues. For example, we study the complex energy level spacing ratio (CLSR) [50, 51] instead of the conventional real energy level spacing ratio (RLSR) [9]. Additionally, we calculate a suitably defined OTOC for non-Hermitian systems. [52]. Recently, the time evolution of the OTOC has been studied for the PTKR model and its classical counterpart in the vicinity of a $\mathcal{PT}$ symmetry-breaking transition in parameter space.[53, 52, 54]. Our study extends this study to include the entire phase diagram of the PTKR model. It provides critical insights into the interplay of the $\mathcal{PT}$ symmetry-breaking transition and the transition from integrability to chaos.
The rest of the paper is structured as follows. In section 2, we introduce the PTKR model. In section 3, we discuss the two diagnostics of quantum chaos – complex level spacing ratio and out-of-ordered correlator. The main results based on these diagnostics are presented in section 4. Finally, we conclude in section 5 discussing our work’s implications and future directions.
2 Model
@@
$\mathcal{P}\hat{\theta}\mathcal{P}^{-1}=-\hat{\theta}\,\quad\quad\mathcal{P}\hat{p}\mathcal{P}^{-1}=-\hat{p}$ (3)
@@
3 Methods
3.1 Energy Level Spacing Ratio
@@
$\xi_{K}=\dfrac{z_{K}^{\text{NN}}-z_{K}}{z_{K}^{\text{NNN}}-z_{K}}=r_{K}\exp{i\theta_{K}}$ ) (7)
The complex level spacing ratio (CLSR) can be defined in two ways: One is by averaging over the magnitudes of $\xi$ to obtain $\langle r\rangle$ (derived from the set of $r_{K}$ values), and the other is averaging over the angular distribution to obtain $-\langle\cos{\theta}\rangle$ ) ⟩ (derived from the set of $\theta_{K}$ values). Only $\langle r\rangle$ can be defined in the limit when the spectrum becomes real. However, in this limit, Eq. 7 does not yield the standard real level spacing ratio (RLSR) for a real spectrum. We, henceforth, refer only to the mean value $\langle r\rangle$ as the complex level spacing ratio (CLSR). Calculations of $-\langle\cos{\theta}\rangle$ ) ⟩ can be found in the supplementary material. The universality classes for random matrices that describe Hermitian matrices in terms of Gaussian Ensembles have suitable generalization to complex matrices [49]. We calculate the CLSR for these universality classes, one of which the PTKR model belongs.
There are two computational techniques to improve the statistics in our calculation. The first is to replace $m$ in Eq. (1) by $m+\Delta m_{p}$, where $\Delta m_{p}$ is a small random number selected independently for each $p$. This ensures that no underlying symmetries leading to unwanted degeneracies occur. This is important as degeneracies must be avoided when calculating level spacing ratios. The second is to average over a small window of values of the kicking strength $K$. This is done to avoid any accidental resonance in the $K$ values. Although, this averaging wasn’t performed while plotting the colormaps to keep the transitions intact.
3.2 Out-of-time-order correlator
@@
Figure 1: Left The CLSR and RLSR as a function of the kicking strength $K$, calculated for $N=8005$ with $\hbar=0.2$. Horizontal lines represent analytical values of RLSR for Poisson and GOE statistics. The figure shows that the CLSR is as good an indicator of the transition from integrability to chaos as the more commonly employed RLSR. Right The CLSR and RLSR as a function of the non-Hermiticity parameter $\lambda$, calculated for $N=4095$ with $\hbar=0.2$ and $K=0.15$. Horizontal lines represent the values of CLSR for the Poisson and GOE distributions. The figure resembles the transition observed for the CLSR on the left got by varying $K$ with $\lambda=0$. However, it is obtained by tuning $\lambda$ instead of $K$. This suggests that the system can transition from a $\mathcal{PT}$-symmetric integrable phase to a $\mathcal{PT}$-symmetric chaotic phase with increasing non-Hermiticity while the value of $K$ is kept constant.
@@
We choose $p_{0}=\hbar_{\text{eff}}k_{0}\in[-\pi,\pi]$ and $\sigma=4$. Numerically, $|\Psi\rangle$ is represented in a finite basis of eigenstates $|k\rangle$, $k\in[-N;N-1]$. The Hamiltonian is expressed as a matrix in this basis of $|k\rangle$ states, and all calculations are also performed w.r.t this basis. In both, the calculation of the OTOC and CLSR, an “effective value” of the Planck constant $\hbar_{\text{eff}}$ has been employed as a parameter to tune the amount of discreetness of the momentum operator [42]. This allows one to make a connection with the classical model by taking the limit $\hbar_{\text{eff}}\rightarrow 0$.
@@
Figure 2: Left: The CLSR for varying values of the kicking strength $K$ and the non-Hermiticity $\lambda$, computed for $N=6005$, $\hbar=0.2$. The dotted lines T1 and T2 represents the transitions across the $\mathcal{PT}$-symmetric integrable phase to $\mathcal{PT}$ symmetry broken chaotic phase and the $\mathcal{PT}$-symmetric chaotic phase to $\mathcal{PT}$ symmetry broken chaotic phase, respectively. Line T* is the Hermitian transition from the integrable to chaotic phase. T3 marks a transition similar to T* in which $K$ is held constant and $\lambda$ varied. Center: The CLSR across the transition Main panel: T1 Inset T2, calculated for $N=4095$. The horizontal lines show the standard values of the CLSR for the GOE and Poisson distributions. Right: The absolute value of the maximum imaginary part of an energy eigenvalue, $\alpha$, across the transition Main panel: T1 Inset T2, calculated for $N=4095$. It can be seen that while $\alpha$ shows an abrupt change along T2, it seems to increase smoothly along T1. The CLSR on the other hand, shows an abrupt transition along both T1 and T2.
4 Results
We now present the results of the calculations discussed in Sec. 3 for the PTKR model.
4.1 CLSR
@@
For the Hermitian case ($\lambda=0$), the mean RLSR can be used to identify the transition from integrability to chaos. In what follows, we show that the CLSR can also be used for the same purpose and takes on characteristic values in the two regimes.
@@
In Fig. 1, we show the mean RLSR as well as the CLSR with varying values of the kicking strength $K$, for $\hbar=0.2$ and system-size $N=8005$. We notice that the transition points are reasonably independent of the value of $N$. Thus, in what follows, the system size is chosen to be large enough so that all quantities are well converged. We find that both the CLSR and RLSR display a transition at the same value of $K$. However, the values in the two regimes are different. Furthermore, the values of the CLSR and RLSR that are obtained by the PTKR model in the chaotic limit of the Hermitian case match with those obtained from averaging over purely random matrices in the Gaussian Orthogonal Ensemble (GOE). The transition in the Hermitian case has been highlighted in the color plot as T* for aid in comparison.
@@
We now consider the non-Hermitian case ($\lambda\neq 0$), which is of particular interest in this study. In Fig. 2, we show the behavior of the CLSR for varying values of kicking strength $K$ and the non-Hermiticity $\lambda$, computed for $\hbar_{\text{eff}}=0.2$, and system-size $N=6005$. We define $\alpha=\text{max}\{\text{Im}(E)\}$ which only becomes non-zero when the spectrum starts to possess imaginary eigenvalues of the energy or, equivalently, when $\mathcal{PT}$-symmetry breaks (see supplementary material). We find that the phase diagram derived from the CLSR consists of three regimes: a $\mathcal{PT}$-symmetric integrable phase, a $\mathcal{PT}$-symmetric chaotic phase, and $\mathcal{PT}$-symmetry broken chaotic phase.
@@
There are three possible transitions that can occur: 1) From the $\mathcal{PT}$-symmetric integrable phase to the $\mathcal{PT}$-symmetry broken chaotic phase, which we label $T_{1}$, 2) from the $\mathcal{PT}$-symmetric chaotic phase to the $\mathcal{PT}$-symmetry broken chaotic phase, which we label $T_{2}$ and 3) from the $\mathcal{PT}$-symmetric integrable phase to $\mathcal{PT}$-symmetric chaotic phase, which we label $T_{3}$. These are shown in Fig. 2. $T_{3}$ is similar to the transition seen in the Hermitian case. This is because, along $T_{3}$, we see that the CLSR goes from 0.50 to 0.57, which is exactly what is observed when going from the integrable to the chaotic regime for $\lambda=0$. For transition $T_{2}$, we see a corresponding abrupt change in the value of $\alpha$ across the transition from the $\mathcal{PT}$-symmetric integrable regime to the $\mathcal{PT}$-symmetry broken chaotic regime. It can be seen that $\mathcal{PT}$-symmetry breaking at the transition $T_{1}$ does not occur abruptly as can be expected for a system with finite Hilbert space dimension (See Fig. 2). Therefore, the exact position of the transition point cannot be determined with great precision; however, the CLSR does show a clear transition. We define a threshold value of $\alpha$ above which the $\mathcal{PT}$-symmetry is assumed to be broken. The threshold, keeping in mind numerical errors, is taken to be $\alpha\leq 10^{-10}$ for unbroken $\mathcal{PT}$-symmetry. We observe that the CLSR takes on the same value of 0.57 in the $\mathcal{PT}$-symmetric chaotic regime as it does in the chaotic regime of the Hermitian model. However, we find that in the $\mathcal{PT}$-symmetric chaotic regime, the value of $\alpha$ can be greater than the threshold of $10^{-10}$, demonstrating that the highest imaginary value of the eigenenergies is not a particularly good diagnostic to detect the absence of $\mathcal{PT}$-symmetry.
@@
It is interesting to note that the transition from the $\mathcal{PT}$-symmetric integrable phase to the $\mathcal{PT}$-symmetric chaotic also occurs across $T_{3}$. The hermitian case $\lambda=0$ exhibits a transition from $\mathcal{PT}$- symmetric integrable to the $\mathcal{PT}$-symmetric chaotic phase as indicated by $T*$. A similar transition can thus be expected when one varies $K$ for $\lambda$ sufficiently small but not equal to zero as seen in the left panel of Fig. 2. The same transition can also be effected by varying $\lambda$ but not $K$ as indicated by $T_{3}$. This is somewhat surprising as it seems to lack any Hermitian counterpart. One is thus able to induce chaos, as seen from the measured CLSR values, by increasing $\lambda$ whilst holding $K$ to the traditional integrable value. The nature of this transition is intriguing and requires further study, which we defer to future work.
@@
Figure 3: The OTOC $C(t)$ and state norm $\mathcal{N}(t)$ vs $t$ across the transition along Left $T_{1}$ and Right $T_{2}$ shown in Fig. 2, calculated for system size $N=2^{14}$. For T1 (Left), the values $(\lambda,K)=\{(0.001,0.1),(0.002,0.456),(0.003,1.135)\}$ correspond to the $\mathcal{PT}$-symmetry unbroken integrable phase, while the values $(\lambda,K)=\{(0.006,5.734),(0.01,15.0)\}$ correspond to the $\mathcal{PT}$-symmetry broken chaotic phase. For T2 (right), the values $(\lambda,K)=\{(10^{-5},30),(10^{-4},30),(10^{-3},30)\}$ correspond to the $\mathcal{PT}$-symmetry unbroken chaotic phase, while the values $(\lambda,K)=\{(0.005,30),(0.001,30.0)\}$ correspond to the $\mathcal{PT}$-symmetry broken chaotic phase. In both cases, the $\mathcal{PT}$-symmetric broken phase displays a late-time exponential growth of the OTOC and the state norm.
4.2 OTOC